A 10-foot board is to be cut into 2 pieces whose lengths are in the ratio 4:1. Find the number of feet in the length of the shorter piece.
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(48558)

A 10-foot board is to be cut into 2 pieces whose lengths are in the ratio 4:1. Find the number of feet in the length of the shorter piece.
------------
4:1 is the same as 4x:x
-------------------------------
Equation:
4x + x = 10
5x = 10
x = 2 ft. (the shorter piece)
4x = 8 ft.(the longer piece)
==========================
Cheers,
Stan H.

Answer by checkley77(12569)

4+1=5
10/5*1=2*1=2 ft. is the length of the shorter piece.
10/5*4=2*4=8 ft. is the length of the longer piece.
Proof:
2+8=10
2:8=1:4